Wednesday May 25
The day shift began with the Antiproton Source (Pbar) stack at 103.5 E10, and the Tevatron (TeV) in shot setup.
Operations established store 4175 at 10:02 AM with an initial luminosity of 62.12E30.
Anti Proton Source (Pbar) resumed stacking, beam to MCenter, and NuMI.
Operations put the Recycler into shot setup at 12:58 PM to stash antiprotons.
Pbar returns to stacking at 2:17 PM.
At 2:24 PM the Pbar rings are in access to repair 2 water leaks.
At 5:15 PM the Pbar water leaks are repaired and they are staking once again.
Thursday May 26
The midnight shift began with Operations monitoring store 4175, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.
Store 4175 was terminated at 9:02 AM.
At 9:15 AM, SY120 and NuMI turned off for accesses.
Shot Setup to Tev started at 9:20 AM.
Operations established store 4176 at 11:34 AM with an initial luminosity of 86.35E30.
Pbar and MCenter resumes taking beam.
At 12:30 PM Main Injector (MI), Booster, Pre-Vault, MiniBoone, and Meson were turned off for access to do necessary repairs.
At 4:41 PM beam returns to all areas.
At 8:47 PM Booster RF station #12 tripped off, it continued tripping and was bypassed at 10:30 PM so that Booster could resume transporting beam.
Shot setup to the Recycler began at 10:36 PM and was completed by 11:26 PM.
Friday May 27
The midnight shift began with Operations monitoring store 4176, with Pbar stacking, a RR stash of 43.3E10,and with MiniBooNE, NuMI, and MCenter taking beam.
The Plan for Friday and the Weekend
The plan for today is to terminate the store at 11:00 AM, D0 will make a 1 hour access, followed by Tev studies and then put the TeV into shot setup at 2:00 PM. The plans for the weekend are to stack, stash, and store.
Accelerator Update Archive
For Tevatron luminosity charts and the current status of Fermilab's
accelerators and detectors (live!), please go to Fermilab Now
Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: email@example.com.